# Krylov exponents and power spectra for maximal quantum chaos: an EFT approach

### v2

@@ 3 Simple shift-symmetric correlators and their Krylov exponents
In this section, we focus on analysing four classes of correlation functions of the operator $\hat{V}(t)$ that solve the constraints discussed in subsection 2.3 via their Lanczos coefficients to test the UOGH and the proposed inequality (25). The primary question we address here is whether the Lanczos coefficients are asymptotically linear, in line with the UOGH (17), and whether the bound (25) is satisfied, which, in the EFT, since the Lyapunov exponent is maximal, requires $\lambda_{K}=2\pi T$.
@@ 3.1 First order shift symmetry constraints: the Schwarzian correlator
An example of a theory for which the shift symmetry constraint equations are first order is Schwarzian quantum mechanics, for which the couplings between the hydro mode and the matter fields are $c_{01}=\Delta_{V}$ and $c_{10}=1$ Blake:2017ris , giving $f_{0}(\lambda)=\lambda\Delta_{V}$ and $f_{1}(\lambda)=1$, see (27), and, for these $f_{i}(\lambda)$, the solution to eq. (41) is the Schwarzian correlator
@@
For large $n$, these coefficients $b_{n}$ are asymptotically linear in $n$: $b_{n}\sim\frac{\pi}{\beta}n$. Note that the asymptotic slope $\alpha$ is $\pi T$ and independent of $f_{0}$ and $f_{1}$. Also, these coefficients are a smooth function of $n$, so the asymptotic slope of $b_{n}$ determines the Krylov exponent, $\lambda_{K}=2\pi T$. This is the only possible value of $\lambda_{K}$ consistent with the bound (25) and $\lambda_{L}=2\pi T$, as required by the EFT. Hence, the solutions to eqs. (40) and (41) always give a $\lambda_{K}$ consistent with eq. (25), for any $f_{0}$ and $f_{1}$.
@@ 3.2 Second order shift symmetry constraints
up to a normalisation factor [^3] Note that this correlation function $C(t)$ is not $\beta$-periodic in imaginary time since only time-ordered correlators are required to be periodic in thermal time. See also the massive scalar field example from Avdoshkin:2022xuw for an explicit $C(t)$ that is not periodic in thermal time. . Eq. (47) is analytic in the neighbourhood of $t=0$, and, for $0<\kappa<1$, is analytic within the strip $|\text{Im}(t)|<2\pi/\lambda=\beta$, which is twice as wide as the thermal strip. Using the Fourier transform identity in eq. (135) while restricting to $0<\kappa<1$, the associated power spectrum reads
@@
where we have set $\lambda=2\pi/\beta$, as required by consistency in the EFT Blake:2021wqj . In the large $n$ limit, the Lanczos coefficients behave as
@@
Hence, in this case, we find a half-maximal Krylov exponent, $\lambda_{K}=\pi T=\frac{1}{2}\lambda_{L}$. Note that the Lanczos coefficients in eq. (49) asymptote very quickly to a special case within the family of Lanczos coefficients associated with continuous Hahn polynomials derived in Gamayun:2025hvu . As shown there, for these coefficients, the Krylov complexity grows exponentially, with a growth rate, in the conventions used here, equal to $\lambda_{K}=\frac{1}{2}\lambda_{L}$. A simple numerical analysis for eq. (49) indeed confirms this. We will defer an extensive discussion of the observation that this particular example leads to submaximal chaos to subsection 3.5.
@@
We can calculate the Lanczos coefficients using the Toda-method detailed in Appendix A and, in this particular case, since the spectrum has only a single frequency, the Krylov space is two-dimensional and there are only two Lanczos coefficients viswanath1994recursion . [^4] Every (non-vanishing) frequency determines two Lanczos coefficients. Hence, $L$-frequencies lead to at most a $2L$-dimensional Krylov space viswanath1994recursion . As a result, the Krylov exponent is not well-defined, and the UOGH is not applicable.
@@ 3.3.3 Simple examples
Identifying, at large $n$, the $b_{n}\sim(\lambda_{K}/2)n$, this leads to the Krylov exponent $\lambda_{K}=\lambda_{L}=2\pi T$.
@@ 3.4 Even and odd-order shift symmetry constraints
Note that similar discrete power spectra were also considered in Anegawa:2024wov ; Anegawa:2024yia .
@@ 3.5 Interpretation of the submaximal Krylov exponents within the EFT
In the previous sections, we considered several examples of two-point functions that satisfy the shift symmetry constraints, as well as a minimal set of physical constraints, see Table 1. Within this space of solutions, a subset satisfying a periodicity condition in imaginary frequency was simple enough that we could calculate the Lanczos coefficients analytically and find $\lambda_{K}=\lambda_{L}/2$. This result is in tension with the conjectured bound $\lambda_{L}\leq\lambda_{K}\leq 2\pi T$. To critically assess these results, it is important to contrast the regime of validity and assumptions of the EFT of maximal chaos and the UOGH.
  • Infinite vs finite $N$. The Krylov exponents obtained in this work arise from the autocorrelator of the bare field, which corresponds to the leading-order contribution in the $1/N$ expansion to the autocorrelator of the full field; see eq. (9). This level of approximation is sufficient to reproduce the maximal Lyapunov exponent in the OTOC, and as demonstrated in Blake:2017ris ; Blake:2021wqj , where the addition of the shift symmetry constraints ensures the expected exponential growth of OTOCs governed by the Lyapunov exponent. However, we find that in the leading-order truncation, some correlation functions yield half-maximal Krylov exponents.
@@
    Finally, let us reiterate that the generalised bound [^7] This refers to the bound even when assuming monotonicity of the Lanczos coefficients, whether formulated in terms of the decay rate of the power spectrum or the slope of the Lanczos coefficients themselves. in eq. (3) remains conjectural. Partial evidence in support of this conjecture has been presented, notably by Avdoshkin:2019trj . More recently, Gu:2021xaj showed that the bound holds in a broad class of theories. However, the derivation in Gu:2021xaj crucially relies on a precise factorisation structure of the OTOC at all orders in the $1/N$ expansion. While the EFT framework employed here does exhibit such factorisation at leading order in $1/N$, it imposes no constraints on higher-order corrections, and moreover, there are technical obstacles to imposing factorisation at higher order in $1/N$, leaving the validity of the provided argument in this context an open question.
@@
  • IR vs. UV degrees of freedom. An important question is what mechanism ensures $\lambda_{K}=2\pi T$ in maximally chaotic systems. One proposed route is via the analytic structure of the two-point function, which is controlled, for example, by conformal symmetry or a UV regulator. For instance, in QFTs, the thermal Wightman correlator defines a well-regulated inner product for local operators, and in 2D CFTs on a plane, it takes a universal form fixed entirely by symmetry. Local continuum QFTs have poles in $C(t)$ when the operators are coincident on the thermal circle, at $t=\pm i\beta/2$, and when the Lanczos coefficients are smooth functions of $n$, this leads to $\lambda_{K}=2\pi T$ Dymarsky:2021bjq . This behaviour appears even in integrable or free theories, where the exponential growth of Krylov complexity reflects the correlator’s analytic structure rather than genuine chaos. We refer the reader to Rabinovici:2025otw for a recent comprehensive discussion.
@@
    To isolate chaos-sensitive contributions, some have proposed adding a UV cutoff or discretising the theory on a lattice Camargo:2022rnt ; Dymarsky:2021bjq ; Avdoshkin:2022xuw . In our case, the EFT naturally includes a UV cutoff since all the UV degrees of freedom are integrated out. Yet we find that the Krylov bound $\lambda_{K}=2\pi T$ is not automatically realised. [^8] Note that if one naively takes the infinite temperature limit in these examples, an apparent contradiction arises with a proven bound Parker:2018yvk stating that, at infinite temperature, the Krylov exponent must bound the Lyapunov exponent from above. However, as the temperature approaches infinity, the IR/UV scale separation required to define the local effective action in the EFT breaks down because the length scale of thermal fluctuations is smaller than the UV cutoff scale. Consequently, the finite-temperature EFT examples considered here do not straightforwardly extend to an infinite-temperature theory, which necessarily involves all degrees of freedom beyond the EFT description. While the shift symmetry is sufficient to ensure maximal Lyapunov growth, we have found that it is not sufficient to ensure maximal Krylov exponent.
@@ 4.1 General solution in the frequency domain
matching, as expected Blake:2017ris , the conformal correlator.
@@ 4.2 KMS-symmetric solutions
The general solution in the frequency domain obtained in eq. (83) represents a large class of solutions, even for a fixed set of roots $\alpha_{i}$, since one can choose any periodic function $G$, as long as the resulting power spectrum obeys the physical constraints and is Fourier-invertible. We will focus on three particular choices for $G$ and see, for one of these choices, that once the physical conditions are imposed, a familiar structure emerges: the power spectrum is closely related to the set of quasi-normal modes (QNM) predicted from the thermal product formula of Dodelson:2023vrw .
@@ 4.3 Relation to the Thermal Product Formula
The general expression in eq. (113) was conjectured by positing a minimal set of constraints on the power spectrum, motivated by thermal conditions and assuming conformal symmetry and holography. In particular, besides the general properties of unitarity and KMS, Dodelson:2023vrw demanded that the power spectrum 1) only display singularities at isolated simple poles in the complex plane, 2) not possess any zeroes (i.e $1/\Phi(\omega)$ is an entire function), and 3) a constraint on the asymptotic behaviour requiring that along any ray in the complex plane that asymptotically avoids poles, the power spectrum satisfies
@@
Earlier, in eq. (112), we found solutions to the shift symmetry constraints that bear a strong resemblance to the thermal product formula (113) of Dodelson:2023vrw . There are, however, several important differences:
  • In the construction of Dodelson:2023vrw , the two-sided thermal correlator is obtained by solving a wave equation in a black hole background, leading to a power spectrum where all singularities are isolated and simple poles. These poles correspond to QNMs of the black hole geometry and encode the real-time decay of thermal perturbations. The assumption of simplicity is therefore not an auxiliary condition but intrinsic to the physical origin of the correlator.
@@
  • In Dodelson:2023vrw , the power spectrum did not possess any zeros in the complex $\omega$-plane, implying that $1/\Phi(\omega)$ is an entire function. This non-vanishing property arises from solving the bulk wave equation in a smooth black hole background, where the ingoing boundary condition at the horizon and the analytic structure of the scattering potential ensure a zero-free correlator Dodelson:2023vrw . In contrast, our (holography-inspired) construction derives the power spectrum from general thermal principles together with maximal chaos in the out-of-time-order correlator. As seen, for example, in eq. (98), power spectra derived from the shift symmetry constraints that satisfy a minimal set of physical conditions can vanish at $\omega=0$.
@@ 5 Discussion
In this work, we contrasted prominent diagnostics of quantum chaos, the OTOC and Krylov space-based probes like Krylov complexity and the UOGH, in an EFT describing maximally chaotic quantum systems. In the EFT construction, the exponential growth of OTOCs with the maximal Lyapunov exponent arises due to a shift symmetry, which also imposes a constraint on the two-point functions of simple operators. In parallel, it is conjectured, through the bound (25), that all maximally chaotic systems have $\lambda_{K}=2\pi T$. We tested this conjecture in the EFT framework by considering various two-point functions that satisfy the shift symmetry constraints as well as a minimal set of physical constraints appropriate for thermal systems, including KMS symmetry, unitarity, and reality from Hermiticity of operators. We computed the corresponding Lanczos coefficients, and, within the set of two-point functions for which we could analytically calculate the Lanczos coefficients, we did indeed find some with $\lambda_{K}=2\pi T$, but also some with $\lambda_{K}=\pi T$
Next, we analysed the general structure of solutions to the shift symmetry equation after imposing the minimal set of physical constraints. In the time domain, the correlator solutions are Meijer G-functions. In the frequency domain, a subset of our solutions resemble the “thermal product formula” of Dodelson:2023vrw , sharing the same overall structure but differing in important analytic features: the holographic power spectra have only simple poles - no poles of order $n>1$ - and no zeroes, whereas power spectra in the EFT allow for more general singularity structures, including both higher-order poles and zeroes.
@@
While we were able to solve the finite $m$ shift symmetry constraint equation in generality, we were only able to analytically calculate the Lanczos coefficients and Krylov exponents for a subset. Interestingly, for all correlators in that subset, the Lanczos coefficients $b_{n}$ are asymptotically linear and thus satisfy the UOGH. However, the full set of solutions consistent with shift symmetry is vast, and we cannot exclude the possibility of more exotic behaviours. Further work on this, both analytical and numerical, is warranted. A natural next step would be to determine what analytic structure in the two-point function gives rise to non-smooth $b_{n}$, including staggering, and to determine whether such solutions are physically realised within the constraints of the EFT. Early explorations in this direction can be found in Camargo:2022rnt ; Dymarsky:2021bjq ; Avdoshkin:2022xuw .
@@
Our analysis is at leading order in $1/N$, and there is currently no formulation of the EFT that captures subleading corrections due to difficulties in systematically implementing KMS symmetry beyond leading order Crossley:2015evo ; Glorioso:2017fpd . Should such a formulation become tractable, it would enable a more refined analysis of thermal correlators and potentially sharpen the EFT predictions for Krylov space-based quantities and their comparison to the UOGH.
@@
  • Other EFT frameworks, such as the effective descriptions developed in Haehl:2018izb , where conformal symmetry plays a central role. Understanding whether our findings map onto theirs, and what distinguishes truly universal chaotic behaviour from model-specific features, could clarify the extent to which the SYK model, with its Schwarzian and large $q$ limits, is a typical or exceptional chaotic theory.
  • Our focus is on one-dimensional systems, and extending these ideas to higher dimensions is an important challenge. The construction of a systematic EFT for operator dynamics in higher-dimensional chaotic systems is still lacking, though recent progress Knysh:2024asf has opened new possibilities.
  • We have restricted our attention to maximally chaotic systems, but a recent proposal Gao:2023wun has extended the EFT approach to systems with submaximal chaos. Developing tools to probe Krylov exponents in this class of thermalising quantum systems is another interesting next step.
@@
Within the general set of solutions, given in eq. (83), for a particular choice of periodic factor $G$ inspired by the conformal limit of SYK, we found power spectra that closely resemble the thermal product formula introduced in Dodelson:2023vrw , and further developed in Bhattacharya:2025vyi . In these examples, this structure arises from the structure of the general solution, which follows from the shift symmetry constraints and physical requirements, and our choice of $G$, without assuming a specific holographic dual.
We also identified several other families of solutions within the general construction that violate the assumptions of having no zeroes and simple poles underlying the thermal product formula of Dodelson:2023vrw . This provides an interesting contrast: our approach is based solely on imposing maximal chaos and the minimal conditions required for a consistent thermal system, rather than the stricter assumptions used in the thermal product derivation. While no explicit holographic dual is assumed, the effective field theory framework is nonetheless strongly guided and motivated by holographic intuition.
Our approach parallels several recent developments in thermal CFTs Iliesiu:2018fao ; Marchetto:2023xap ; Barrat:2025wbi ; Barrat:2025nvu and holographic correlators Buric:2025anb , where thermal two-point functions are derived from minimal analytic, OPE, and symmetry input, supplemented by KMS invariance. Remarkably, such basic ingredients suffice to reconstruct non-trivial thermal observables. Exploring how far methods based on extracting thermal systems from consistency conditions can be pushed, and clarifying their limitations, is key to understanding the interplay between universal features and model‑dependent aspects of thermal physics.

### v3

@@ 3 Simple shift-symmetric correlators and their Krylov exponents
In this section, we focus on analysing four classes of correlation functions of the operator $\hat{V}(t)$ that solve the constraints discussed in subsection 2.3 via their Lanczos coefficients to test the UOGH and the proposed inequality (25). The primary question we address here is whether the Lanczos coefficients are asymptotically linear, in line with the UOGH (17), and whether the shift symmetry constraints ensure that the Krylov exponent is maximal $\lambda_{K}=2\pi T$.
@@ 3.1 First order shift symmetry constraints: the Schwarzian correlator
An example of a theory for which the shift symmetry constraint equations are first order is Schwarzian quantum mechanics, for which the couplings between the hydro mode and the matter fields are $c_{01}=\Delta_{V}$ and $c_{10}=1$ Blake et al. (2018), giving $f_{0}(\lambda)=\lambda\Delta_{V}$ and $f_{1}(\lambda)=1$, see (27), and, for these $f_{i}(\lambda)$, the solution to eq. (41) is the Schwarzian correlator
@@
For large $n$, these coefficients $b_{n}$ are asymptotically linear in $n$: $b_{n}\sim\frac{\pi}{\beta}n$. Note that the asymptotic slope $\alpha$ is $\pi T$ and independent of $f_{0}$ and $f_{1}$. Also, these coefficients are a smooth function of $n$, so the asymptotic slope of $b_{n}$ determines the Krylov exponent, $\lambda_{K}=2\pi T$. This value of $\lambda_{K}$ is maximal as expected for maximally chaotic systems. Hence, the solutions to eqs. (40) and (41) always give a $\lambda_{K}$ consistent with eq. (25), for any $f_{0}$ and $f_{1}$.
@@ 3.2 Second order shift symmetry constraints
up to a normalisation factor [^5] Note that this correlation function $C(t)$ is not $\beta$-periodic in imaginary time since only time-ordered correlators are required to be periodic in thermal time. See also the massive scalar field example from Avdoshkin et al. (2022) for an explicit $C(t)$ that is not periodic in thermal time. . Eq. (47) is analytic in the neighbourhood of $t=0$, and, for $0<\kappa<1$, is analytic within the strip $|\text{Im}(t)|<2\pi/\lambda=\beta$, which is twice as wide as the thermal strip. Using the Fourier transform identity in eq. (135) while restricting to $0<\kappa<1$, the associated power spectrum reads
@@
where we have set $\lambda=2\pi/\beta$, as required by consistency in the EFT Blake and Liu (2021). In the large $n$ limit, the Lanczos coefficients behave as
@@
Hence, in this case, we find a half-maximal Krylov exponent, $\lambda_{K}=\pi T$. Note that the Lanczos coefficients in eq. (49) asymptote very quickly to a special case within the family of Lanczos coefficients associated with continuous Hahn polynomials derived in Gamayun et al. (2025). As shown there, for these coefficients, the Krylov complexity grows exponentially, with a growth rate, in the conventions used here, equal to $\lambda_{K}=\pi T$. A simple numerical analysis for eq. (49) indeed confirms this. We will defer an extensive discussion of the observation that this particular example leads to submaximal chaos to subsection 3.5.
@@
We can calculate the Lanczos coefficients using the Toda-method detailed in Appendix A and, in this particular case, since the spectrum has only a single frequency, the Krylov space is two-dimensional and there are only two Lanczos coefficients Viswanath and Müller (1994). [^6] Every (non-vanishing) frequency determines two Lanczos coefficients. Hence, $L$-frequencies lead to at most a $2L$-dimensional Krylov space Viswanath and Müller (1994). As a result, the Krylov exponent is not well-defined, and the UOGH is not applicable.
@@ 3.3.3 Simple examples
Identifying, at large $n$, the $b_{n}\sim(\lambda_{K}/2)n$, this leads to the Krylov exponent $\lambda_{K}=2\pi T$.
@@ 3.4 Even and odd-order shift symmetry constraints
Note that similar discrete power spectra were also considered in Anegawa et al. (2024); Anegawa and Watanabe (2024).
@@ 3.5 Interpretation of the submaximal Krylov exponents within the EFT
In the previous sections, we considered several examples of two-point functions that satisfy the shift symmetry constraints, as well as a minimal set of physical constraints, see Table 1. Within this space of solutions, a subset satisfying a periodicity condition in imaginary frequency was simple enough that we could calculate the Lanczos coefficients analytically and find $\lambda_{K}=\pi T$. This result indicates that within the EFT of maximal chaos, the shift symmetry guarantees maximal Lyapunov growth in the OTOC diagnostic as defined in the EFT framework, while it does not uniquely determine the Krylov exponent extracted from thermal two-point functions. Interpreting this in terms of the conjectured bound $\lambda_{L}\leq\lambda_{K}\leq 2\pi T$ requires care, since at finite temperature the exponents depend on the choice of regularization scheme. The apparent tension arises when one adopts a broad interpretation of the conjecture, in which $\lambda_{L}$ is identified with the Lyapunov exponent extracted from the OTOC as defined in Blake et al. (2018); Blake and Liu (2021), and/or when one assumes that maximal chaos as seen through the OTOC is a physical statement and thus independent of this choice. Let us highlight that for maximally chaotic systems such as those described by the EFT construction, $\lambda_{L}$ is expected to be universal ($2\pi T$) and robust against changes in regularization. See footnote 4 for references and discussion. To critically assess these results, it is also important to contrast the regime of validity and assumptions of the EFT of maximal chaos and the UOGH.
  • Infinite vs finite $N$. The Krylov exponents obtained in this work arise from the autocorrelator of the bare field, which corresponds to the leading-order contribution in the $1/N$ expansion to the autocorrelator of the full field; see eq. (9). This level of approximation is sufficient to reproduce the maximal Lyapunov exponent in the OTOC, and as demonstrated in Blake et al. (2018); Blake and Liu (2021), where the addition of the shift symmetry constraints ensures the expected exponential growth of OTOCs governed by the Lyapunov exponent. However, we find that in the leading-order truncation, some correlation functions yield half-maximal Krylov exponents.
    This leading-order feature is most extreme in the oscillatory solution of eq. (52). Here, the operator is effectively confined to a finite-dimensional Krylov subspace at leading order, yet the OTOC remains maximally chaotic. This apparent paradox is due to the structure of the large-$N$ expansion. The maximal Lyapunov growth in the OTOC arises from the exchange of the hydrodynamic mode, which is a $1/N$ effect. In contrast, the operator growth diagnosed by the leading-order $O(1)$ contribution to the correlation function remains blind to the chaotic nature of the EFT mediated by the hydrodynamic mode.
@@
    Finally, let us reiterate that the generalised bound [^9] This refers to the bound even when assuming monotonicity of the Lanczos coefficients, whether formulated in terms of the decay rate of the power spectrum or the slope of the Lanczos coefficients themselves. in eq. (3) remains conjectural. Partial evidence in support of this conjecture has been presented, notably by Avdoshkin and Dymarsky (2020). More recently, Gu et al. (2022) showed that the bound holds in a broad class of theories. However, the derivation in Gu et al. (2022) crucially relies on a precise factorisation structure of the OTOC at all orders in the $1/N$ expansion. While the EFT framework employed here does exhibit such factorisation at leading order in $1/N$, it imposes no constraints on higher-order corrections, and moreover, there are technical obstacles to imposing factorisation at higher order in $1/N$, leaving the validity of the provided argument in this context an open question.
@@
  • IR vs. UV degrees of freedom. An important question is what mechanism ensures $\lambda_{K}=2\pi T$ in maximally chaotic systems. One proposed route is via the analytic structure of the two-point function, which is controlled, for example, by conformal symmetry or a UV regulator. For instance, in QFTs, the thermal Wightman correlator defines a well-regulated inner product for local operators, and in 2D CFTs on a plane, it takes a universal form fixed entirely by symmetry. Local continuum QFTs have poles in $C(t)$ when the operators are coincident on the thermal circle, at $t=\pm i\beta/2$, and when the Lanczos coefficients are smooth functions of $n$, this leads to $\lambda_{K}=2\pi T$ Dymarsky and Smolkin (2021). This behaviour appears even in integrable or free theories, where the exponential growth of Krylov complexity reflects the correlator’s analytic structure rather than genuine chaos. We refer the reader to Rabinovici et al. (2025) for a recent comprehensive discussion.
@@
    To isolate chaos-sensitive contributions, some have proposed adding a UV cutoff or discretising the theory on a lattice Camargo et al. (2023); Dymarsky and Smolkin (2021); Avdoshkin et al. (2022). In our case, the EFT naturally includes a UV cutoff since all the UV degrees of freedom are integrated out. Yet we find that the Krylov bound $\lambda_{K}=2\pi T$ is not automatically realised. [^10] Note that if one naively takes the infinite temperature limit in these examples, an apparent contradiction arises with a proven bound Parker et al. (2019) stating that, at infinite temperature, the Krylov exponent must bound the Lyapunov exponent from above. However, as the temperature approaches infinity, the IR/UV scale separation required to define the local effective action in the EFT breaks down because the length scale of thermal fluctuations is smaller than the UV cutoff scale. Consequently, the finite-temperature EFT examples considered here do not straightforwardly extend to an infinite-temperature theory, which necessarily involves all degrees of freedom beyond the EFT description. While the shift symmetry is sufficient to ensure maximal Lyapunov growth, we have found that it is not sufficient to ensure maximal Krylov exponent.
@@ 4.1 General solution in the frequency domain
matching, as expected Blake et al. (2018), the conformal correlator.
@@ 4.2 KMS-symmetric solutions
The general solution in the frequency domain obtained in eq. (83) represents a large class of solutions, even for a fixed set of roots $\alpha_{i}$, since one can choose any periodic function $G$, as long as the resulting power spectrum obeys the physical constraints and is Fourier-invertible. We will focus on three particular choices for $G$ and see, for one of these choices, that once the physical conditions are imposed, a familiar structure emerges: the power spectrum is closely related to the set of quasi-normal modes (QNM) predicted from the thermal product formula of Dodelson et al. (2024).
@@ 4.3 Relation to the Thermal Product Formula
The general expression in eq. (113) was conjectured by positing a minimal set of constraints on the power spectrum, motivated by thermal conditions and assuming conformal symmetry and holography. In particular, besides the general properties of unitarity and KMS, Dodelson et al. (2024) demanded that the power spectrum 1) only display singularities at isolated simple poles in the complex plane, 2) not possess any zeroes (i.e $1/\Phi(\omega)$ is an entire function), and 3) a constraint on the asymptotic behaviour requiring that along any ray in the complex plane that asymptotically avoids poles, the power spectrum satisfies
@@
Earlier, in eq. (112), we found solutions to the shift symmetry constraints that bear a strong resemblance to the thermal product formula (113) of Dodelson et al. (2024). There are, however, several important differences:
  • In the construction of Dodelson et al. (2024), the two-sided thermal correlator is obtained by solving a wave equation in a black hole background, leading to a power spectrum where all singularities are isolated and simple poles. These poles correspond to QNMs of the black hole geometry and encode the real-time decay of thermal perturbations. The assumption of simplicity is therefore not an auxiliary condition but intrinsic to the physical origin of the correlator.
@@
  • In Dodelson et al. (2024), the power spectrum did not possess any zeros in the complex $\omega$-plane, implying that $1/\Phi(\omega)$ is an entire function. This non-vanishing property arises from solving the bulk wave equation in a smooth black hole background, where the ingoing boundary condition at the horizon and the analytic structure of the scattering potential ensure a zero-free correlator Dodelson et al. (2024). In contrast, our (holography-inspired) construction derives the power spectrum from general thermal principles together with maximal chaos in the out-of-time-order correlator. As seen, for example, in eq. (98), power spectra derived from the shift symmetry constraints that satisfy a minimal set of physical conditions can vanish at $\omega=0$.
@@ 5 Discussion
In this work, we contrasted prominent diagnostics of quantum chaos, the OTOC and Krylov space-based probes like Krylov complexity and the UOGH, in an EFT describing maximally chaotic quantum systems. In the EFT construction, the exponential growth of OTOCs with the maximal Lyapunov exponent arises due to a shift symmetry Blake et al. (2018); Blake and Liu (2021), which also imposes a constraint on the two-point functions of simple operators. A natural question is then whether this same structure is strong enough to enforce maximal Krylov growth. Motivated by the conjectured relation (25), which suggests that $\lambda_{K}$ should be bounded by the chaos scale and saturates $2\pi T$ in maximally chaotic systems, we explored this question within the EFT. To this end, we constructed families of thermal two-point functions consistent with the shift symmetry constraints as well as a minimal set of physical constraints appropriate for thermal systems, including KMS symmetry, unitarity, and reality from Hermiticity of operators. We computed the corresponding Lanczos coefficients, and, within the set of two-point functions for which we could analytically calculate the Lanczos coefficients, we did indeed find some with $\lambda_{K}=2\pi T$, but also some with $\lambda_{K}=\pi T$
Next, we analysed the general structure of solutions to the shift symmetry equation after imposing the minimal set of physical constraints. In the time domain, the correlator solutions are Meijer G-functions. In the frequency domain, a subset of our solutions resemble the “thermal product formula” of Dodelson et al. (2024), sharing the same overall structure but differing in important analytic features: the holographic power spectra have only simple poles - no poles of order $n>1$ - and no zeroes, whereas power spectra in the EFT allow for more general singularity structures, including both higher-order poles and zeroes.
@@
While we were able to solve the finite $m$ shift symmetry constraint equation in generality, we were only able to analytically calculate the Lanczos coefficients and Krylov exponents for a subset. Interestingly, for all correlators in that subset, the Lanczos coefficients $b_{n}$ are asymptotically linear and thus satisfy the UOGH. However, the full set of solutions consistent with shift symmetry is vast, and we cannot exclude the possibility of more exotic behaviours. Further work on this, both analytical and numerical, is warranted. A natural next step would be to determine what analytic structure in the two-point function gives rise to non-smooth $b_{n}$, including staggering, and to determine whether such solutions are physically realised within the constraints of the EFT. Early explorations in this direction can be found in Camargo et al. (2023); Dymarsky and Smolkin (2021); Avdoshkin et al. (2022).
@@
Our analysis is at leading order in $1/N$, and there is currently no formulation of the EFT that captures subleading corrections due to difficulties in systematically implementing KMS symmetry beyond leading order Crossley et al. (2017); Glorioso et al. (2017). Should such a formulation become tractable, it would enable a more refined analysis of thermal correlators and potentially sharpen the EFT predictions for Krylov space-based quantities and their comparison to the UOGH.
@@
  • Other EFT frameworks, such as the effective descriptions developed in Haehl and Rozali (2018), where conformal symmetry plays a central role. Understanding whether our findings map onto theirs, and what distinguishes truly universal chaotic behaviour from model-specific features, could clarify the extent to which the SYK model, with its Schwarzian and large $q$ limits, is a typical or exceptional chaotic theory.
  • Our focus is on one-dimensional systems, and extending these ideas to higher dimensions is an important challenge. The construction of a systematic EFT for operator dynamics in higher-dimensional chaotic systems is still lacking, though recent progress Knysh et al. (2024) has opened new possibilities.
  • We have restricted our attention to maximally chaotic systems, but a recent proposal Gao and Liu (2023) has extended the EFT approach to systems with submaximal chaos. Developing tools to probe Krylov exponents in this class of thermalising quantum systems is another interesting next step.
@@
Within the general set of solutions, given in eq. (83), for a particular choice of periodic factor $G$ inspired by the conformal limit of SYK, we found power spectra that closely resemble the thermal product formula introduced in Dodelson et al. (2024), and further developed in Bhattacharya et al. (2025). In these examples, this structure arises from the structure of the general solution, which follows from the shift symmetry constraints and physical requirements, and our choice of $G$, without assuming a specific holographic dual.
We also identified several other families of solutions within the general construction that violate the assumptions of having no zeroes and simple poles underlying the thermal product formula of Dodelson et al. (2024). This provides an interesting contrast: our approach is based solely on imposing maximal chaos and the minimal conditions required for a consistent thermal system, rather than the stricter assumptions used in the thermal product derivation. While no explicit holographic dual is assumed, the effective field theory framework is nonetheless strongly guided and motivated by holographic intuition.
Our approach parallels several recent developments in thermal CFTs Iliesiu et al. (2018); Marchetto et al. (2024); Barrat et al. (2025b, a) and holographic correlators Burić et al. (2025), where thermal two-point functions are derived from minimal analytic, OPE, and symmetry input, supplemented by KMS invariance. Remarkably, such basic ingredients suffice to reconstruct non-trivial thermal observables. Exploring how far methods based on extracting thermal systems from consistency conditions can be pushed, and clarifying their limitations, is key to understanding the interplay between universal features and model‑dependent aspects of thermal physics.